# Point Cloud Synthesis Using Inner Product Transforms

### v4

@@ 3.2 The IP-Encoder
As an alternative to the machine learning model above, we also investigate whether the inversion of an IPT can be turned into an optimisation problem. Due to the differentiability of our approximation $\widehat{\mathrm{IPT}}$ and given a known IPT, minimising a loss function, such as the one described in Lemma 1, using backpropagation should result in a suitable approximation of the unknown point cloud one wishes to recover. However, while this works in theory, it requires the input and target to be a ‘correct’ $\widehat{\mathrm{IPT}}$ of a point cloud. In practice, this is not guaranteed: Generative models may sample from their latent spaces and output samples that are close to being an $\widehat{\mathrm{IPT}}$ of a point cloud without actually satisfying all structural constraints. To some extent, we argue that this is even the desired behaviour since a generative model would otherwise just rehash its inputs [16]. Hence, any optimisation-based method will not necessarily result in realistic reconstructions in a generative setting. We nevertheless investigate such an optimisation procedure in an ablation study in the appendix (Figure S.5).
@@ Results.
Our reconstructions (Table 1 and Figure 2(a)) are of high quality, despite a comparatively simple architecture. Our IP-Encoder consistently ranks second in terms of the EMD, known to be the most suitable metric to evaluate reconstruction quality [44]. Notably, our loss term does not use the EMD. Our method also exhibits substantially reduced training times compared to other methods: The IP-Encoder model requires approximately $30\text{\,}\mathrm{min}$ on a single GPU, compared to $192\text{\,}\mathrm{h}$ for PointFlow and more than of $550\text{\,}\mathrm{h}$ for LION. We additionally perform an ablation study concerning the loss term, showing that the combination of CD and IPT is crucial for high reconstruction quality. As Table 2 shows, a joint loss yields the best quality. In line with prior work, CD on its own tends to adversely affect point-cloud density [2], resulting in larger (i.e. worse) EMD scores. Conversely, training only with a ECT loss shows that the reconstructed follow the density of the underlying object, but may become blurry. This is partially due to the low resolution of $64$, making the combination of CD (for details) and the ECT (for global density and structure) crucial. While it would also possible to use the EMD as a loss term, we find that our IPT-64 loss function is substantially faster than the EMD (0.000 $00\,64\text{\,}\mathrm{s}$ versus $0.029\,13\text{\,}\mathrm{s}$), making it the preferred loss function.
@@
Table 4 reports numerical results, whereas Figure 2(b) depicts generated samples from each class. We include different quality metrics and note that, despite its conceptual simplicity, our model consistently ranks among the best two models in terms of MMD-EMD, and exhibits performance on a par with more elaborate architectures. This is noteworthy since the IP-VAE is not constrained to produce exact IPTs, hence its outputs may contain artefacts. The fact that our IP-Encoder model (remaining fixed throughout this experiment) can reconstruct point clouds from such out-of-sample data without additional regularisation or retraining, shows that the IPT describes a stable and expressive latent space. We observe fast training times, with the IP-VAE taking approximately $15\text{\,}\mathrm{min}$, making the full training pipeline run in less than $1\text{\,}\mathrm{h}$. This is in stark contrast to models like LION ($550\text{\,}\mathrm{h}$) or SoftFlow ($\approx$144\text{\,}\mathrm{h}$\$). Finally Table 3 shows that our inference times on commodity hardware are orders of magnitude faster than existing models.
@@
Table 5 depicts numerical scores (cf. Figure S.3 in the appendix for sample point clouds). Unlike uniform subsampling, our downsampling model preserves uniform density of each object, showing that it has learned the underlying shape. We observe some loss in quality compared to reconstruction without downsampling, which is to be expected. Despite potential compounding errors, the loss in quality remains low. Due to the similar setup, we can compare results with Table 1 and we observe that even when downsampling, we still perform on a par with the best model in the reconstruction task. Notably, the IP-Encoder has not been retrained for this task, so the IPT of the subsampled point cloud is out-of-distribution. This indicates that our model has captured the ‘true’ underlying shape characteristics.
@@
Table 6 and Figure S.5 in the appendix depict the numerical results and show some reconstructed samples, respectively. Comparing the result with Table 1, we observe that backpropagation outperforms the IP-Encoder model. Table S.1 in the appendix provides the full set of results and an ablation with respect to the scale and resolution. We observe that all approximations become gradually coarser as the resolution is decreased. As expected, we observe problems with vanishing gradients for large values of $\lambda$, since the sigmoid function approaches an indicator function in this case. This requires balancing $\lambda$ with respect to the resolution, and we empirically observed that around a quarter of the IPT resolution is sufficient for high-quality results, which is how we pick the parameter in practice. It is important to note that while this method of inverting works for an IPT without any artifacts, it fails to accurately converge if the IPT is only an approximation of a true IPT, as is the case with a generated IPT. This also further motivates the use of a learned method for reconstruction.
@@ 4.5 Interpolation in Latent Space
As our final experiment, we consider the linear interpolation between the IPT of two point clouds. In the setting of infinite directions, Theorem 2 guarantees that we only encounter valid IPTs along each interpolation. However, in practice, given our use of image-generative models, it is not guaranteed that our input is an IPT in the strict sense. Hence, it is crucial to understand the characteristic properties of the IPT in practice since not every sum of sigmoid functions is guaranteed to correspond to a proper shape (in other words, IPTs are not surjective). In our discretised representation, interpolating between IPTs can be achieved by interpolating per pixel. To visually assess the quality of the latent space, we reconstruct each intermediary IPT during the interpolation using the IP-Encoder. Figure 3 depicts the resulting point clouds; we observe that the intermediary point clouds still remain plausible reconstructions. Overall, this serves to highlight the utility of the IPT latent space.
@@ 5 Discussion
An unfortunate limitation arising from the disconnect between theory and practice is that our injectivity results do not guarantee that one can find suitable directions, in particular given the complex interplay between directions and resolutions of the IPT. We believe that additional guidance for selecting such parameters would be helpful. The current formulation of the IPT is also not invariant with respect to rotations, which could be a desired property for some applications (or, at the very least, might make the method more robust in the regime of smaller sample sizes). However, with increasing sample size, we observe that data augmentation enables IPT-based models to ‘learn’ equivariance despite not being intrinsically invariant (see Figure S.7 in the appendix for preliminary results). Moreover, if point clouds become high-dimensional, our direction-sampling procedure becomes inefficient—this is not a problem for the shape-generation tasks we are tackling in this paper, though. A larger limitation in point-cloud generation involves the question of evaluation metrics: In the absence of ground-truth information, it is hard to choose the ‘best’ model (unlike point-cloud reconstruction, where well-defined metrics exist). In terms of other generative-model evaluation metrics like COV and 1-NNA, our models exhibits only middling performance, ranking slightly better in terms of COV than in terms of 1-NNA. This is in stark contrast to its excellent reconstruction performance and its strong MMD scores. Taken together, this indicates that the $\widehat{\mathrm{IPT}}$ is capable of generating high-quality point clouds, but its generative part, based on simple VAEs, is not sufficiently powerful to provide diverse representations.
@@ Acknowledgments
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The core contributions of our work are clearly stated and reflected in the abstract and presented in the introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: An extensive discussion is provided in a separate section as well as potential avenues of future work.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate ”Limitations” section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We provide clear statements as to what our assumptions are for all theoretical claims. Full proofs are provided in the appendix for brevity.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide detailed information for the models used in the experiments and an interested reviewer should be able to recreate our experiments based on the manuscript. Moreover, the full implementation is publicly available at https://github.com/aidos-lab/inner-product-transforms.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide the code with adequate instruction for reproducing all our experiments in the supplementary materials. Our implementation and full experimental setup are publicly available.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provide full details.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Where applicable we provide error bars, if in the literature it is not common to report error bars for certain metrics we follow the standards and not provide them.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer ”Yes” if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Extensive tables are provided and we list both training and inference times for all our models. The hardware specifications of the compute resources are also provided in the main text.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We fully comply with the code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We provide an impact statement.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We do not provide pretrained models and / or scraped data.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All datasets and models used are properly referenced.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We will provide an anonymized version of the code with instructions for running the main experiments during the review process and after the review process the code will become publicly available.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Our data and experiments do not use data obtained from human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We do not use data obtained from crowdsourcing nor data from human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: No LLM has been used for the writing, coding, editing nor the research.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v5

@@ 3.2 The IP-Encoder
As an alternative to the machine learning model above, we also investigate whether the inversion of an IPT can be turned into an optimisation problem. Due to the differentiability of our approximation $\widehat{\mathrm{IPT}}$ and given a known IPT, minimising a loss function, such as the one described in Lemma 1, using backpropagation should result in a suitable approximation of the unknown point cloud one wishes to recover. However, while this works in theory, it requires the input and target to be a ‘correct’ $\widehat{\mathrm{IPT}}$ of a point cloud. In practice, this is not guaranteed: Generative models may sample from their latent spaces and output samples that are close to being an $\widehat{\mathrm{IPT}}$ of a point cloud without actually satisfying all structural constraints. To some extent, we argue that this is even the desired behaviour since a generative model would otherwise just rehash its inputs [16]. Hence, any optimisation-based method will not necessarily result in realistic reconstructions in a generative setting. We nevertheless investigate such an optimisation procedure in an ablation study in the appendix (Fig. S.5).
@@ Results.
Our reconstructions (Table 1 and Fig. 2(a)) are of high quality, despite a comparatively simple architecture. Our IP-Encoder consistently ranks second in terms of the EMD, known to be the most suitable metric to evaluate reconstruction quality [44]. Notably, our loss term does not use the EMD. Our method also exhibits substantially reduced training times compared to other methods: The IP-Encoder model requires approximately $30\text{\,}\mathrm{min}$ on a single GPU, compared to $192\text{\,}\mathrm{h}$ for PointFlow and more than of $550\text{\,}\mathrm{h}$ for LION. We additionally perform an ablation study concerning the loss term, showing that the combination of CD and IPT is crucial for high reconstruction quality. As Table 2 shows, a joint loss yields the best quality. In line with prior work, CD on its own tends to adversely affect point-cloud density [2], resulting in larger (i.e. worse) EMD scores. Conversely, training only with a ECT loss shows that the reconstructed follow the density of the underlying object, but may become blurry. This is partially due to the low resolution of $64$, making the combination of CD (for details) and the ECT (for global density and structure) crucial. While it would also possible to use the EMD as a loss term, we find that our IPT-64 loss function is substantially faster than the EMD (0.000 $00\,64\text{\,}\mathrm{s}$ versus $0.029\,13\text{\,}\mathrm{s}$), making it the preferred loss function.
@@
Table 4 reports numerical results, whereas Fig. 2(b) depicts generated samples from each class. We include different quality metrics and note that, despite its conceptual simplicity, our model consistently ranks among the best two models in terms of MMD-EMD, and exhibits performance on a par with more elaborate architectures. This is noteworthy since the IP-VAE is not constrained to produce exact IPTs, hence its outputs may contain artefacts. The fact that our IP-Encoder model (remaining fixed throughout this experiment) can reconstruct point clouds from such out-of-sample data without additional regularisation or retraining, shows that the IPT describes a stable and expressive latent space. We observe fast training times, with the IP-VAE taking approximately $15\text{\,}\mathrm{min}$, making the full training pipeline run in less than $1\text{\,}\mathrm{h}$. This is in stark contrast to models like LION ($550\text{\,}\mathrm{h}$) or SoftFlow ($\approx$144\text{\,}\mathrm{h}$\$). Finally Table 3 shows that our inference times on commodity hardware are orders of magnitude faster than existing models.
@@
Table 5 depicts numerical scores (cf. Fig. S.3 in the appendix for sample point clouds). Unlike uniform subsampling, our downsampling model preserves uniform density of each object, showing that it has learned the underlying shape. We observe some loss in quality compared to reconstruction without downsampling, which is to be expected. Despite potential compounding errors, the loss in quality remains low. Due to the similar setup, we can compare results with Table 1 and we observe that even when downsampling, we still perform on a par with the best model in the reconstruction task. Notably, the IP-Encoder has not been retrained for this task, so the IPT of the subsampled point cloud is out-of-distribution. This indicates that our model has captured the ‘true’ underlying shape characteristics.
@@
Table 6 and Fig. S.5 in the appendix depict the numerical results and show some reconstructed samples, respectively. Comparing the result with Table 1, we observe that backpropagation outperforms the IP-Encoder model. Table S.1 in the appendix provides the full set of results and an ablation with respect to the scale and resolution. We observe that all approximations become gradually coarser as the resolution is decreased. As expected, we observe problems with vanishing gradients for large values of $\lambda$, since the sigmoid function approaches an indicator function in this case. This requires balancing $\lambda$ with respect to the resolution, and we empirically observed that around a quarter of the IPT resolution is sufficient for high-quality results, which is how we pick the parameter in practice. It is important to note that while this method of inverting works for an IPT without any artifacts, it fails to accurately converge if the IPT is only an approximation of a true IPT, as is the case with a generated IPT. This also further motivates the use of a learned method for reconstruction.
@@ 4.5 Interpolation in Latent Space
As our final experiment, we consider the linear interpolation between the IPT of two point clouds. In the setting of infinite directions, Theorem 2 guarantees that we only encounter valid IPTs along each interpolation. However, in practice, given our use of image-generative models, it is not guaranteed that our input is an IPT in the strict sense. Hence, it is crucial to understand the characteristic properties of the IPT in practice since not every sum of sigmoid functions is guaranteed to correspond to a proper shape (in other words, IPTs are not surjective). In our discretised representation, interpolating between IPTs can be achieved by interpolating per pixel. To visually assess the quality of the latent space, we reconstruct each intermediary IPT during the interpolation using the IP-Encoder. Fig. 3 depicts the resulting point clouds; we observe that the intermediary point clouds still remain plausible reconstructions. Overall, this serves to highlight the utility of the IPT latent space.
@@ 5 Discussion
An unfortunate limitation arising from the disconnect between theory and practice is that our injectivity results do not guarantee that one can find suitable directions, in particular given the complex interplay between directions and resolutions of the IPT. We believe that additional guidance for selecting such parameters would be helpful. The current formulation of the IPT is also not invariant with respect to rotations, which could be a desired property for some applications (or, at the very least, might make the method more robust in the regime of smaller sample sizes). However, with increasing sample size, we observe that data augmentation enables IPT-based models to ‘learn’ equivariance despite not being intrinsically invariant (see Fig. S.7 in the appendix for preliminary results). Moreover, if point clouds become high-dimensional, our direction-sampling procedure becomes inefficient—this is not a problem for the shape-generation tasks we are tackling in this paper, though. A larger limitation in point-cloud generation involves the question of evaluation metrics: In the absence of ground-truth information, it is hard to choose the ‘best’ model (unlike point-cloud reconstruction, where well-defined metrics exist). In terms of other generative-model evaluation metrics like COV and 1-NNA, our models exhibits only middling performance, ranking slightly better in terms of COV than in terms of 1-NNA. This is in stark contrast to its excellent reconstruction performance and its strong MMD scores. Taken together, this indicates that the $\widehat{\mathrm{IPT}}$ is capable of generating high-quality point clouds, but its generative part, based on simple VAEs, is not sufficiently powerful to provide diverse representations.